OpenStudy (anonymous):

Find all polar coordinates of point P = (9, 75°). ?

OpenStudy (anonymous):

P = (9, (75+360k)°), for all integers k P = (−9, (255+360k)°), for all integers k

OpenStudy (anonymous):

When you add 180 degrees, terminal point will be directly opposite from original position, but if you then let r = -9, you will be back to original position: P = (−9, (75+180)°) = (−9, 255°) When you add or subtract a multiple of 360 degrees to an angle, terminal point will be in same location.